# Robust Distributed Gradient Descent to Corruption over Noisy Channels

### v1

@@ Appendix D Proof of Lemma 3
$H_{1}L_{1}=\eta_{1}L(\theta_{1})+\eta_{1}\langle\tilde{g}_{1},\theta_{2}-\theta_{1}\rangle+\Big{(}\frac{\eta_{1}\alpha}{2}+\frac{\alpha_{0}}{2}\Big{)}\|\theta_{2}-\theta_{1}\|_{2}^{2}\\ -\frac{\mu_{0}}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}-\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{1}\rangle.$
@@
$\leqslant\frac{\alpha_{0}}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle$
$\qquad\quad-\eta_{1}\langle\nabla L(\theta_{1}),\theta_{2}-\theta_{1}\rangle-\Big{(}\frac{\eta_{1}\alpha}{2}+\frac{\alpha_{0}}{2}\Big{)}\Big{\|}\theta_{2}-\theta_{1}\|_{2}^{2}$
$\leqslant\frac{\alpha_{0}}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}^{2}M^{2}}{2(\eta_{1}\alpha\!+\!\alpha_{0})}$
$=\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}M^{2}}{4\alpha}$
@@ Appendix E Proof of Lemma 4
$=\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
@@ Appendix F Proof of Lemma 5
$\geqslant\eta_{t}L(\theta_{t})+\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+\frac{(H_{t}+\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
@@
$=\eta_{t}L(\theta_{t})\!+\!\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\!+\!\frac{(H_{t}\!+\!\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
@@
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})\\ -\eta_{t}L(\theta_{t})-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{(H_{t}\!+\!\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$ (21)
@@
$\eta_{t}(L(\theta_{t+1})-L(\theta_{t}))-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\\ -\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}+\eta_{t}(L(\theta_{t})-L(\theta_{t+1}))-\!\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$
Combining the fact that $\frac{\eta_{t}}{H_{t}}\leqslant\frac{\alpha}{M}$ and the $M$-smoothness of $L$, we obtain
$\eta_{t}\left(L(\theta_{t+1})\!-\!L(\theta_{t})\!-\!\langle\nabla L(\theta_{t}),\theta_{t+1}\!-\!\theta_{t}\rangle\right)\!-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant\frac{\eta_{t}M}{2}\|\theta_{t+1}\!-\!\theta_{t}\|_{2}^{2}-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}=0.$ (22)
By the convexity of $L$ and the Cauchy–Schwartz inequality,
$\eta_{t}(L(\theta_{t})-L(\theta_{t+1}))-\!\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant\frac{\eta_{t}M^{2}}{2\alpha}.$ (23)
Combining (21), (F) and (F), we conclude that
$H_{t}G_{t}\leqslant H_{t-1}G_{t-1}+\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle+\frac{\eta_{t}M^{2}}{2\alpha}$
@@ Appendix G Proof of Corollary 1
$G_{t}=\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}\sum_{k=1}^{t}\eta_{k}}{2\alpha H_{t}}\\ +\frac{\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle}{H_{t}}$
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}\\ +\frac{R_{\Theta}C(t)}{m}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
@@ Appendix H Proof of Lemma 6
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}\\ +R_{\Theta}t^{r}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{\alpha R_{\Theta}^{2}}{2}e^{-\frac{\alpha}{M}(t-1)}+\frac{M^{2}}{2\alpha}\\ +R_{\Theta}t^{r}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (24)
@@ Appendix I Proof of Theorem 3
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}\prod_{k=2}^{t}(1-\frac{\eta_{k}}{H_{k}})\Big{)}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}\\ +R_{\Theta}t^{r}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$

### v2

@@ Appendix D Proof of Lemma 3
$H_{1}L_{1}=\eta_{1}L(\theta_{1})+\eta_{1}\langle\tilde{g}_{1},\theta_{2}-\theta_{1}\rangle+\eta_{1}\alpha\|\theta_{2}-\theta_{1}\|_{2}^{2}\\ -\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}-\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{1}\rangle.$
@@
$\leqslant\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle$
$\qquad\quad-\eta_{1}\langle\nabla L(\theta_{1}),\theta_{2}-\theta_{1}\rangle-\eta_{1}\alpha\|\theta_{2}-\theta_{1}\|_{2}^{2}$
$\leqslant\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle$
$\quad\quad\quad+\eta_{1}K\|\theta_{2}-\theta_{1}\|_{2}-\eta_{1}\alpha\|\theta_{2}-\theta_{1}\|_{2}^{2}$
$\leqslant\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}K^{2}}{4\alpha}$
@@ Appendix E Proof of Lemma 4
$\geqslant\frac{H_{t-1}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$
Hence, we have
$h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})$ $\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle$ (23)
$\qquad+\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$ (24)
@@ Appendix F Proof of Lemma 5
$\geqslant\eta_{t}L(\theta_{t})+\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
@@
$=\eta_{t}L(\theta_{t})\!+\!\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\!+\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
@@
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})\\ -\eta_{t}L(\theta_{t})-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$ (25)
@@
$-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$
Then, we have
$-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant\eta_{t}K\|\theta_{t+1}-\theta_{t}\|_{2}-\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant\frac{\eta_{t}^{2}K^{2}}{2H_{t}\alpha}\leqslant\frac{\eta_{t}K^{2}}{2\alpha},$ (26)
where the last inequality from the fact that $H_{t}\geqslant\eta_{t}$.
Combining (25) and (F), we conclude that
$H_{t}G_{t}\leqslant H_{t-1}G_{t-1}+\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle+\frac{\eta_{t}K^{2}}{2\alpha}$
@@ Appendix G Proof of Corollary 1
$G_{t}=\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{K^{2}\sum_{k=1}^{t}\eta_{k}}{2\alpha H_{t}}\\ +\frac{\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle}{H_{t}}$
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{K^{2}}{2\alpha}\\ +\frac{R_{\Theta}C(t)}{m}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
@@ Appendix H Proof of Lemma 6
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{K^{2}}{2\alpha}\\ +R_{\Theta}t^{r}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{\alpha R_{\Theta}^{2}}{2}e^{-\frac{\alpha}{M}(t-1)}+\frac{K^{2}}{2\alpha}\\ +R_{\Theta}t^{r}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (27)
@@ Appendix I Proof of Theorem 3
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}\prod_{k=2}^{t}(1-\frac{\eta_{k}}{H_{k}})\Big{)}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{K^{2}}{2\alpha}\\ +R_{\Theta}t^{r}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$